The lengths of the sides of a right triangle are three consecutive integers. Write and solve an equation to find three integers.
Answer by JThomson(12)


Let the sides be x, (x+1) and (x+2), where (x+2) is the hypotenuse;

(x+2)^2 = (x+1)^2 + x^2

x^2+4x+4 = x^2+2x+1+x^2

x^2-2x-3=0

(x-3)(x+1)=0

x = 3 since x must be positive

(x+1)=4 and (x+2)=5

Hence the sides are 3, 4 and 5 respectively